my question is this
solve the quadratic equation using the formula
1/1+x -1/3-x = 6/35

Answer by checkley75(3666)

1/1+X-1/3-X=6/35
X-X=6/35-1+1/3
0=6/35-1+1/3
NO SOLUTION!
ARE YOU SURE YOU COPIED THE EQUATION CORRECTLY?